A MEASURE OF DISCRIMINATION BETWEEN TWO RESIDUAL LIFE-TIME DISTRIBUTIONS AND ITS APPLICATIONS

NADER EBRAHIMI^{1} AND S. N. U. A. KIRMANI^{2}

1Division of Statistics and Statistical Consulting Laboratory, Northern Illinois University, DeKalb, IL 60115-2854, U.S.A.

2Department of Mathematics, University of Northern Iowa, Cedar Falls, IA 50614, U.S.A.

(Received March 20, 1995)

Abstract. A measure of discrepancy between two residual-life distributions is proposed on the basis of Kullback-Leibler discrimination information. Properties of this measure are studied and the minimum discrimination principle is applied to obtain the proportional hazards model.

Key words and phrases: Hazard function, cumulative hazard function, mean residual lifetime function, NBU, IFR, Kullback-Leibler discrimination information, MDI principle.